Suppose that Gopal goes to the stargazing event for exactly one hour each night on 5 different nights. On 4 of the nights, the average number of shooting star sightings per hour is 2. However, on the fifth night, there is a meteor shower and the average number of sightings is 6 per hour. What is the expected number of total shooting stars that Gopal observes in the five nights of stargazing?
Answer #1

On 4 of the nights, the average number of shooting star sightings per hour is 2.

Gopal goes to the stargazing event for exactly one hour each night so the total number of shooting stars he observes in those 4 nights is: {(2*1)+(2*1)+(2*1)+(2*1)} =2+2+2+2=8

On the 5th night he observed on an averageof 6 sightings per hour. He was there for 1 hour only. So he ovserved : 6*1=6 shooting stars.

Therefore the expected number of total shooting stars that Gopal observes in the five nights of stargazing , is : (8+6)=14 . (ans)